OpenStudy (anonymous):

You roll a standard number cube. Find P(3 or 4). HELP? URGENT. WILL MEDAL.

OpenStudy (anonymous):

@dan815 Do you think you could possibly help me with this and a few other questions?

OpenStudy (dumbcow):

there are 6 possible numbers .... 3 or 4 are 2 possible numbers out of the 6 ---> 2/6 = 1/3

OpenStudy (anonymous):

Do i multiply those two fractions or is 1/3 the answer?

OpenStudy (anonymous):

oh okay thank you both

OpenStudy (anonymous):

2/6 is your outcome but don't forget to reduce that comes out to 1/3

OpenStudy (anonymous):

You roll a standard number cube. Find P(number greater than 2). For this, I got 2/3 as my final answer..Is this correct?

OpenStudy (anonymous):

4/6 reduce to 2/3

OpenStudy (anonymous):

Yes correct! :)

OpenStudy (anonymous):

Okay, thank you. And the answer to this is 1/6 correct?... You roll a pair of number cubes. What is the probability of getting even numbers on both cubes?

OpenStudy (anonymous):

One cube includes 6 sides even numbers are 2,4 and 6 which is 3/6 so 3/6 +3/6 =6/12 reduce 1/2

OpenStudy (anonymous):

oh..So it's 1/2?

OpenStudy (anonymous):

yes because you reduce it

OpenStudy (anonymous):

h okay thank you

OpenStudy (anonymous):

Simplify the expression. (–2 – c)(–1) I got -2 + c for this.

OpenStudy (anonymous):

I only see rectangles

OpenStudy (anonymous):

Oh well here's what it was..

OpenStudy (anonymous):

you multiplied right?

OpenStudy (anonymous):

No I just kind of mentally tried to solve it

OpenStudy (anonymous):

looks like it could be 2+c im multiplying -2*-1=2 and- c=-1 so -1*-1=1 neg. times a neg.=pos

OpenStudy (anonymous):

@dan815 can you doublecheck all of these? It'd mean the worlddddd

OpenStudy (anonymous):

especially the last one lol

OpenStudy (anonymous):

^^ haha thank you for helping too btw

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

@zepdrix